Crew 238 Sol Summary Report 13Jan2022
Summary Title: A Taste of Fresh Martian Spirulina
Author’s name: Sionade Robinson
Mission Status: Ongoing
Sol Activity Summary: Research work continues. We concluded our
Spirulina competition with a comparison of colour and density of our
individual tubes. Photogra will show the considerable variation, some
thriving and some dead. We compared notes on our various regimen of
care – noting a more consistent approach to temperature, light and
agitation led to optimum results best. This evening Aga harvested 10g
of Spirulina from the litre jar currently incubating in the Green Hab
and we ate it on crackers. It is a high protein foodstuff, bright
green and viscous with only a faint taste. An EVA to Area G Candor
Chasma to extend sampling was successfully completed and an Emergency
Simulation of tunnel breaches executed effectively.
Look Ahead Plan: On our last day at MDRS the crew seekpermission for a
walking EVA for four participants followed by removal of temporary
labyrinth from sand north of Science Dome. Completion of cleaning
programme and packing ready to depart.
Anomalies in work: None.
Weather: Cloudier than yesterday and colder. No wind
Crew Physical Status: Nominal.
EVA: By Rover to Area G.
Reports to be filed: Sol, Operations, EVA Report, Journalist and HSO report.
Support Requested: EVA Request
Additional Support: Wheatgrass harvested (see Operational report) but
not for crew consumption. Left in Airlock for collection for Martian
Dogs who apparently regard it as a treat.